Below are listed those faults for which 2 or more hours were lost.
The cable that guides the ``cable twist'' which conducts the various communication cables from the devices attached to the rotator to the fixed telescope structure broke. This likely occurred because of wear and tear over the last 18 years since its installation. The poor and incomplete documentation seriously complicated identifying the true problem while repairing the cable took relatively long as accessing the cable is rather difficult and it required a certain type of cable of sufficient length.
Apart from having a proper spare cable (which is not expected to be needed any time soon) the only action that was taken is to make a general review of all the status parameters in the TCS to see if they are well defined and can be properly identified (e.g., the parameter for the status of a specific part is connected to a signal that can be traced to that part).
During observations with FIES using the simultaneous ThAr mode it was found that the arc lines were saturated and the target spectrum could barely be seen. The cause was found to be the shutter of the calibration unit in the FIES building not closing. After some time it was found that moving some wires made the shutter work again. During tests later on the problem recurred. No obvious loose contact was found, but after rewiring the connection the problem has not recurred.
Thomas Augusteijn 2009-01-15